# Servo Integrated Nonlinear Model Predictive Control for Overactuated Tiltable-Quadrotors

### v1

@@ Abstract
Quadrotors are widely employed across various domains, yet the conventional type faces limitations due to underactuation, where attitude control is closely tied to positional adjustments. In contrast, quadrotors equipped with tiltable rotors offer overactuation, empowering them to track both position and attitude trajectories. However, the nonlinear dynamics of the drone body and the sluggish response of tilting servos pose challenges for conventional cascade controllers. In this study, we propose a control methodology for tilting-rotor quadrotors based on nonlinear model predictive control (NMPC). Unlike conventional approaches, our method preserves the full dynamics without simplification and utilizes actuator commands directly as control inputs. Notably, we incorporate a first-order servo model within the NMPC framework. Through simulation, we observe that integrating the servo dynamics not only enhances control performance but also accelerates convergence. To assess the efficacy of our approach, we fabricate a tiltable-quadrotor and deploy the algorithm onboard at a frequency of 100Hz. Extensive real-world experiments demonstrate rapid, robust, and smooth pose tracking performance.
@@
Aerial Systems: Mechanics and Control, Motion Control, Overactuation, MPC
@@ I Introduction
Aerial robots have increasingly attracted attentions due to their mobility in three-dimensional space. As the most popular aerial robot, quadrotors have been applied in various areas including transportation, inspection, and search & rescue [1]. Traditional quadrotors are underactuated, meaning that the number of the independent control input (i.e., the rotation speed of four vertically oriented rotors) is fewer than the required six degree-of-freedom (DoF) for full-pose motion, leading to the close relation between attitude and position. However, numerous applications require independent attitude control, particularly in the field of aerial manipulation [2].
@@
Figure 1: Our self-made tiltable-quadrotor is tracking a pose trajectory based on the proposed NMPC method. It demonstrates the capability for independent variation in position and attitude to track different Lemniscate curves.
@@
To overcome the challenges associated with underactuation, researchers have developed two primary ways to achieve fullactuation or overactuation in rotor-based aerial robots. One way involves adding rotors with fixed tilting angles [3, 4], named as fixed-rotor robots. While these robots maintain mechanical and control simplicity comparable to their underactuated counterparts, they encounter issues like constant internal forces, reduced efficiency, and constrained wrench generation in certain directions. Alternatively, the tilt-rotor design introduces servo modules to vary thrust directions, enhancing force distribution and energy efficiency. Moreover, tilt-rotor systems offer the capacity for omnidirectional flight [5], unlocking new potentials in maneuverability and application. Given these advantages, our research focuses on the tiltable-quadrotor, as exemplified in Fig. 1.
@@
The first real-world flight of a quadrotor equipped with tiltable rotors was achieved by Ryll et al., where the control approach is detailed in [6]. Subsequently, the Voliro project [7] introduced a hexacopter with tiltable rotors, marking a pioneering step towards real-world omnidirectional flight. In addition, Senkul et al. [8] developed a quadrotor capable of tilting its rotors along two axes, but they assumed the same thrust generation across all rotors, restricting the vehicle’s maneuverability. Despite these advantages, the flight control for tiltable-multirotor remains challenging.
@@
The control complexity of tiltable-multirotors arises primarily from two factors. First, the introduction of additional DoFs to vary tilting angles significantly increases the system’s nonlinearity. These extra DoFs make the system overactuated, resulting in multiple solutions in control allocation. Second, the inherent dynamics of tilting servos impede prompt motion change for reference tracking. This slow property can be caused by backlash-induced dead time and the servo’s large time constant, which have been reported as critical factor to flight stability [6]. Similar challenges have been observed in other servo-equipped aerial robots, such as SPIDAR [9] and Perching Arm [10].
@@
To address the first challenge, plenty of research have simplified this complex control problem by decoupling into control and allocation components as depicted in Fig. 2a. For control, various strategies based on established control theories have been proposed, including feedback linearization [6], nonlinear inverse dynamics [11], cascade PID [7], LQRI [12], and NMPC [13]. For allocation, Moore–Penrose inverse is typically utilized to map control inputs to specific actuator commands. While these approaches demonstrate efficacy in real-world flights, their cascaded structure may not fully leverage the potential of overactuated systems or elegantly address actuator constraints. To overcome these drawbacks, recent studies begin to integrate control and allocation within a single optimization framework, notably through nonlinear model predictive control (NMPC). This integrated approach has shown promise in conventional quadrotors [14, 15] and is gaining traction for tiltable-multirotors. For instance, Bicego et al. [16] proposed an NMPC framework suitable for various multi-rotor designs, tested on a hexrotor that all motor tilting for the same angle. However, its applicability to drones with independently tiltable rotors remains unverified. Shawky et al. [17] developed an NMPC controller for tiltable hexarotors, albeit only validated in Gazebo simulation. Despite these advancements, the deployment of unified NMPC in real-world tiltable-multirotors is scarcely reported, and thus one highlight in this work is the implementation on a real tiltable-quadrotor.
@@
To address the second challenge, some studies [7, 17] have chosen to overlook the servo effect, which consequently reduce control performance. Ryll et al. [6] bypassed the issue of dead time by using desired instead of measured angles, and implemented a Smith predictor to address the slow servo response. Although effective, this method is complicated as conventional feedback controllers lack predictive property. In contrast, NMPC inherently incorporates prediction, offering an advantageous framework for addressing delayed servo responses. One approach to indirectly consider servomechanism within NMPC is to confine the change rate of resultant wrench, a method proposed by Brunner et al. [13]. While this technique helps manage the servo dynamics, it cannot accurately capture the entire scope of servo behavior. More recent studies [16, 18] integrated the servo model more explicitly by using the derivative of servo angle as a control input. However, this approach, while it moderates the rate of angle change, is unable to comprehensively account for the nuances of dead time and first-order model. Therefore, our work directly consider the servo dynamics inside the NMPC pipeline.
@@
In this work, we introduce a unified nonlinear model predictive control approach for tiltable-quadrotors, as illustrated in 2b. This approach leverages allocation as an external reference, rather than as an module within the control loop. Our method fully incorporates the nonlinear dynamics into the NMPC framework, directly selecting thrust and servo angle as control inputs. Furthermore, we explicitly integrate the servo dynamics within the model. We find that the inclusion of servo dynamics not only improves model fidelity but also facilitates more rapid convergence of the optimizer.
@@
Figure 2: Comparative analysis of the previous and proposed workflows. Unlike the previous cascade structure, the proposed method directly integrates allocation and constraints within the NMPC optimization process. Furthermore, we explicitly model the servo dynamics. This approach has no simplification on the robot model, resulting in the full exploration of the hardware potential.
@@
    We propose an NMPC framework for tiltable-quadrotors, which considers the full dynamics and actuator constraints.
@@
    We explicitly consider the servo dynamics inside the NMPC model, which is verified critical for both performance and convergence.
@@
    The control method is implemented onboard with 100Hz control frequency. As far as the author’s best knowledge, this is the first time an actuator-level NMPC is implemented on real tiltable-quadrotors.
The remainder of the article is organized as follows. The modelling for the tiltable-quadrotor is introduced in Section II. The control approach is presented in Section III, followed by simulation analysis in Section IV. We then show the experimental results in Section V and finally the conclusion in Section VI.
II Modelling
@@ II-B Tiltable-Quadrotor Model
The model for a tiltable-quadrotor consists of three parts. First, we establish the force and torque model generated by rotors. Then we apply this wrench to the rigid-body dynamics to obtain the full robot model. Finally, we introduce a first-order model to explicitly describe the servo dynamics.
II-B1 Rotor Wrench Model
In order to make the established model more general, we assume that $N_{p}$ rotors are fixed on the robot, and we use $d_{i}$ (whose value is ±$\pm 1$) to denote the rotating direction of $i$th motor and ${}^{G}\bm{p}_{ri}$ to denote the rotor’s position in $\left\{\mathcal{G}\right\}$.
Motors rotate their propellers to generate thrust and torque. Here we neglect the angular acceleration of rotors and adopt commonly-used quadratic fit to describe the motor dynamics
@@ II-B1 Rotor Wrench Model
where $\Omega_{i}$ is the rotating speed of the $i$th propeller, ${k}_{t}$ and ${k}_{q}$ are coefficients identified from experiments. It is assumed that the motors are able to achieve this rotation speed $\Omega_{i}$ with negligible transients.
@@
${{}^{R_{i}}\bm{f}_{i}}$ $=\left[0,0,f_{i}\right]^{T},\quad f_{i}\in\left[f_{i,\min},f_{i,\max}\right],$ (2)
${{}^{R_{i}}\bm{\tau}_{i}}$ $=\left[0,0,-d_{i}\ f_{i}\ \frac{k_{q}}{k_{t}}\right]^{T},$ (3)
Based on the coordinate systems, $\alpha_{i}$ influences the rotation from $\left\{\mathcal{R}_{i}\right\}$ to $\left\{\mathcal{E}_{i}\right\}$ along X axis with denotation as
@@
In additional to the established torque, tilting a rotating propeller can generate an additional torque caused by gyroscopic effects. For aerial robot with small size, this effect is 2-3 orders of magnitude smaller [16] and can be treated as high-order terms to neglect. Then the resultant force and torque generated by propellers are derived as
${{}^{G}\bm{f}_{u}}=$ $\sum_{i=1}^{N_{p}}{{}^{G}_{E_{i}}\bm{R}}\ {{}^{E_{i}}_{R_{i}}\bm{R}}\ {{}^{R_{i}}\bm{f}_{i}},$ (5)
$\begin{split}{{}^{G}\bm{\tau}_{u}}=&\sum_{i=1}^{N_{p}}\left({{}^{G}_{E_{i}}\bm{R}}\ {{}^{E_{i}}_{R_{i}}\bm{R}}\ {{}^{R_{i}}\bm{\tau}_{i}}\right.\\ &\left.+^{G}\bm{p}_{ri}\times{{{}^{G}_{E_{i}}\bm{R}}\ {{}^{E_{i}}_{R_{i}}\bm{R}}\ {{}^{R_{i}}\bm{f}_{i}}}\right),\end{split}$ (6)
@@
II-B2 Servo Model
The servo can be modelled with angle, angular velocity or even torque as input. Considering most low-cost servos only support angle control, we hereby model the servo’s motion as a first-order damping model
@@ II-B2 Servo Model
where $\bm{\alpha}=\left[\alpha_{1},\alpha_{2},\dots,\alpha_{N_{p}}\right]^{T}$ denotes the vector including all servo angles, and $\bm{\alpha_{c}}=\left[\alpha_{c1},\alpha_{c2},\dots,\alpha_{cN_{p}}\right]^{T}$ is the symbol for control command.
II-B3 Rigid-Body Model
Neglecting the aerodynamic drag during flying, the motion of robot is caused only by gravitational force and rotor wrench. Using position ${{}^{W}\bm{p}}$, velocity ${}^{W}\bm{v}$, quaternion ${{}^{W}_{G}\bm{q}}$, and angular velocity ${{}^{G}\bm{\omega}}$ (angular velocity of $\left\{\mathcal{G}\right\}$ w.r.t $\left\{\mathcal{W}\right\}$ and expressed in $\left\{\mathcal{G}\right\}$) as states, a six DoF rigid body dynamics can be established as follows
${{}^{W}\dot{\bm{p}}}$ $={{}^{W}\bm{v}},$ (8)
${{}^{W}\dot{\bm{v}}}$ $=\left({{}^{W}_{G}\bm{R}(\bm{q})}\ {{}^{G}\bm{f}_{u}}+{{}^{W}\bm{f}_{d}}\right)/{m}+{{}^{W}\bm{{g}}},$ (9)
${{}^{W}_{G}\dot{\bm{q}}}$ $=\frac{1}{2}\ {{}^{W}_{G}\bm{q}}\circ\mathcal{H}(^{G}\bm{\omega}),$ (10)
${{}^{G}\dot{\bm{\omega}}}$ $=\bm{{I}}^{-1}\left(-{{}^{G}\bm{\omega}}\times\left({\bm{{I}}}\ {{}^{G}\bm{\omega}}\right)+{{}^{G}\bm{\tau}_{u}}+{{}^{G}\bm{\tau}_{d}}\right),$ (11)
where $\bm{R}(\bm{q})$ denotes the rotation matrix converted from quaternion, $\circ$ refers to quaternion multiplication, $\mathcal{H}(\cdot)$ means homogenizing a 3D vector $\mathcal{H}(\bm{p}):=[0,\bm{p}]^{T}$, ${{}^{G}\bm{f}_{u}}$ and ${{}^{G}\bm{\tau}_{u}}$ are expressed in eqs. 5 and 6, ${{}^{W}\bm{f}_{d}}$ and ${{}^{G}\bm{\tau}_{d}}$ are the force and torque caused by disturbances, as well as $m$, ${{}^{W}\bm{{g}}}=[0,0,{-g}]^{T}$, and $\bm{{I}}$ are the mass, the gravity vector, and the inertia matrix, respectively.
The whole model established above are utilized for control and simulation. Note that this model is created for multirotors, and the tiltable-quadrotor is the special case when $N_{p}=4$.

### v2

@@ Abstract
Utilizing a servo to tilt each rotor transforms quadrotors from underactuated to overactuated systems, allowing for independent control of both attitude and position, which provides advantages for aerial manipulation. However, this enhancement also introduces model nonlinearity, sluggish servo response, and limited operational range into the system, posing challenges to dynamic control. In this study, we propose a control approach for tiltable-quadrotors based on nonlinear model predictive control (NMPC). Unlike conventional cascade methods, our approach preserves the full dynamics without simplification. It directly uses rotor thrust and servo angle as control inputs, where their limited working ranges are considered input constraints. Notably, we incorporate a first-order servo model within the NMPC framework. Simulation reveals that integrating the servo dynamics is not only an enhancement to control performance but also a critical factor for optimization convergence. To evaluate the effectiveness of our approach, we fabricate a tiltable-quadrotor and deploy the algorithm onboard at 100 Hz. Extensive real-world experiments demonstrate rapid, robust, and smooth pose-tracking performance.
@@
Aerial Systems: Mechanics and Control, Motion Control, MPC, Overactuated Robots
@@ I Introduction
Aerial robots have increasingly attracted attention due to their mobility in three-dimensional space. As the most popular aerial robot, quadrotors have been applied in various areas including transportation, inspection, and search & rescue [1]. Traditional quadrotors are underactuated, meaning that the number of independent control inputs (i.e., the rotation speed of four vertically oriented rotors) is fewer than the required six degrees of freedom (DoF) for full-pose motion, leading to the independent control of only four flat outputs: three-axis position and yaw angle. However, numerous applications require independent attitude control, particularly in the field of aerial manipulation [2].
@@
Figure 1: Our self-made tiltable-quadrotor is tracking a pose trajectory based on the proposed NMPC method. It demonstrates the capability for independent control in position and attitude to track different Lemniscate curves.
@@
To overcome the challenges associated with underactuation, researchers have developed two primary ways to achieve fullactuation or overactuation in rotor-based aerial robots. One way involves adding rotors with fixed tilting angles [3, 4], named as fixed-rotor robots. While these robots maintain mechanical and control simplicity comparable to their underactuated counterparts, they encounter issues like constant internal forces, reduced efficiency, and constrained wrench generation in certain directions. Alternatively, the tilt-rotor design introduces servo modules to vary thrust directions, reducing internal force and enhancing energy efficiency. If we aim to enable conventional quadrotors to control independently both position and attitude while maintaining the symmetry of roll and pitch, a relatively straightforward idea is to add a servo-tilting mechanism to each rotor, allowing it to rotate along the arm. Therefore, our research focuses on the tiltable-quadrotor, as exemplified in Fig. 1.
@@
The first real-world flight of a quadrotor equipped with tiltable rotors was achieved by Ryll et al., where the control approach is detailed in [5]. Subsequently, the Voliro project [6] introduced a hexacopter with tiltable rotors, marking a pioneering step towards real-world omnidirectional flight. In addition, Senkul et al. [7] developed a quadrotor capable of tilting its rotors along two axes, but they assumed the same thrust generation across all rotors, restricting the vehicle’s maneuverability. Despite these advantages, flight control for tiltable-multirotors remains challenging.
@@
The control complexity of tiltable-multirotors arises primarily from two factors. First, the introduction of additional DoFs to vary tilting angles significantly increases the system’s nonlinearity. These extra DoFs make the system overactuated, resulting in multiple solutions in control allocation. Second, the inherent dynamics of tilting servos impede prompt motion control for agile trajectory tracking. This slow property can be caused by dead time or the servo’s time constant, which has been reported as a critical factor in flight stability [5]. Similar challenges have been observed in other servo-equipped aerial robots, such as SPIDAR [8] and Perching Arm [9].
@@
To address the first challenge, plenty of research has simplified this complex control problem by decoupling it into control and allocation components, as illustrated in Fig. 2a. For control, various strategies have been proposed based on established control theories, including feedback linearization [5], nonlinear inverse dynamics [10], cascade PID [6], LQRI [11], and NMPC [12]. For allocation, Moore–Penrose inverse is typically utilized to map control inputs (for most cases, wrench) to specific actuator commands. Although these approaches demonstrate effectiveness in real-world flights, their cascade structure may not fully leverage the potential of overactuated systems or elegantly address actuators’ delay and constraints. To overcome these drawbacks, recent studies have started to integrate control and allocation within a single optimization framework, notably through nonlinear model predictive control (NMPC). This integrated approach has shown promise in conventional quadrotors [13, 14] and is gaining traction for tiltable-multirotors. For instance, Bicego et al. [15] proposed an NMPC framework suitable for various multi-rotor designs, tested on a hexrotor that all motors can tilt for the same angle. However, its applicability to drones with independently tiltable rotors remains unverified. Shawky et al. [16] developed an NMPC controller for tiltable hexacopters, albeit only validated in Gazebo simulation. Despite these advancements, the deployment of unified NMPC in real-world tiltable-multirotors is scarcely reported, and thus one highlight in this work is the implementation on a real tiltable-quadrotor.
@@
To address the second challenge, some studies [6, 16] have chosen to overlook the servo effect, which consequently reduces the control performance. Ryll et al. [5] implemented a Smith predictor to address the slow servo response. Although effective, this method is complicated since conventional feedback controllers lack the predictive property. In contrast, NMPC inherently incorporates prediction, offering an advantageous framework for systems with sluggish actuators. For example, the works [17], [18] leverage NMPC to control an aerial manipulator comprising both an agile multirotor and relatively slow carriers/winches. For tiltable multirotors, one approach to indirectly consider servomechanism within NMPC is to constrain the change rate of the resultant wrench, a method proposed by Brunner et al. [12]. While this technique helps manage the servo dynamics, it cannot accurately capture the entire scope of servo behavior. More recent studies [15, 19] integrate the actuator model more explicitly by using their derivatives as control inputs (so-called “delta-input formulation” [17]) and trying to constrain the range of these derivatives. Although the “$\Delta u$ formulation” can describe various systems including first-order models, this generality might be unnecessary if the system is already known to be first-order. Furthermore, this way requires an additional integrator to accurately obtain the control command, and the actuators’ measurement noise presents challenges for identifying their derivatives’ boundaries in practice. In contrast, we model the servo as a first-order system, eliminating the need for an integrator and simplifying the identification process.
@@
In this work, we introduce a unified nonlinear model predictive control approach for tiltable-quadrotors, as illustrated in 2b. This approach leverages allocation as an external reference, rather than a module within the control loop. Our method fully incorporates the nonlinear dynamics into the NMPC framework, directly utilizing rotor thrust and servo angle as control inputs. Furthermore, we explicitly integrate the servo as a first-order system within the model. We find that the inclusion of servo dynamics not only improves model fidelity but also facilitates optimization convergence.
@@
Figure 2: Comparative analysis of the previous and proposed workflows. Unlike the previous cascade structure, the proposed method directly integrates constraints and allocation within the NMPC optimization. Furthermore, servos are explicitly modeled. We directly use the servo angle and thrust as commands, requiring no extra integrator as in other research [15].
@@
    We propose a servo-integrated NMPC framework for tiltable-quadrotors, which explicitly considers the servo as a first-order system. This framework has no simplification of the nonlinear robot model, resulting in the full exploration of the hardware potential.
@@
    We point out that due to the servo angle’s nonlinearity and working range, considering the servo model is more critical than thrust for flight performance. In addition, modeling the servo is crucial for optimization convergence. These findings are verified in simulations.
@@
    We thoroughly evaluate the real-world control performance on various references with a self-developed tiltable-quadrotor, where the NMPC controller is running at 100 Hz on an onboard computer. To the best of the authors’ knowledge, this is the first time an actuator-level NMPC is implemented on a real tiltable-quadrotor.
The remainder of the article is organized as follows. The modeling for the tiltable-quadrotor is introduced in Sec. II. The control approach is presented in Sec. III, followed by simulation analysis in Sec. IV. We then show the experimental results in Sec. V and finally the conclusion in Sec. VI.
II Modeling
@@ II-B Tiltable-Quadrotor Model
The model for a tiltable-quadrotor has several parts. First, we establish the rotor model and get its resultant wrench. Then we introduce a first-order model to explicitly describe the servo dynamics. Finally, we apply the wrench to the rigid-body dynamics to obtain the full robot model.
II-B1 Rotor Model
Motors rotate their propellers to generate thrust and torque. Here we neglect the angular acceleration of rotors and adopt commonly used quadratic fit to describe the motor dynamics
@@ II-B1 Rotor Model
where $\Omega_{i}$ is the rotating speed of the $i$th propeller, ${k}_{t}$ and ${k}_{q}$ are coefficients identified from experiments. It is assumed that the motors are able to achieve the rotation speed $\Omega_{i}$ with negligible transients.
II-B2 Resultant Wrench
To make the established model more general, we assume that $N_{p}$ rotors are fixed on the robot, and we use $d_{i}$ (whose value is ±$\pm 1$) to denote the rotating direction of the $i$th motor and ${}^{G}\bm{p}_{r,i}$ to denote its position in $\left\{\mathcal{G}\right\}$.
@@ II-B2 Resultant Wrench
${{}^{R_{i}}\bm{f}_{i}}$ $=\left[0,0,f_{i}\right]^{T},\quad f_{i}\in\left[f_{i,\min},f_{i,\max}\right],$ (2a)
${{}^{R_{i}}\bm{\tau}_{i}}$ $=\left[0,0,-d_{i}\ f_{i}\ \frac{k_{q}}{k_{t}}\right]^{T}.$ (2b)
Based on the coordinate systems, the tilt angle $\alpha_{i}$ influences the rotation from $\left\{\mathcal{R}_{i}\right\}$ to $\left\{\mathcal{E}_{i}\right\}$ along X axis, denoted as
@@
In addition to the established torque, tilting a rotating propeller can generate an additional torque caused by gyroscopic effects. For aerial robots with small sizes, this effect is 2-3 orders of magnitude smaller [15] and can be treated as high-order terms to neglect. Then the resultant force and torque generated by propellers are derived as
${{}^{G}\bm{f}_{u}}=$ $\sum_{i=1}^{N_{p}}{{}^{G}_{E_{i}}\bm{R}}\ {{}^{E_{i}}_{R_{i}}\bm{R}}\ {{}^{R_{i}}\bm{f}_{i}},$ (4a)
$\begin{split}{{}^{G}\bm{\tau}_{u}}=&\sum_{i=1}^{N_{p}}\left({{}^{G}_{E_{i}}\bm{R}}\ {{}^{E_{i}}_{R_{i}}\bm{R}}\ {{}^{R_{i}}\bm{\tau}_{i}}\right.\\ &\left.+^{G}\bm{p}_{r,i}\times{{{}^{G}_{E_{i}}\bm{R}}\ {{}^{E_{i}}_{R_{i}}\bm{R}}\ {{}^{R_{i}}\bm{f}_{i}}}\right),\end{split}$ (4b)
@@
II-B3 Servo Model
The servo can be modeled with angle, angular velocity, or even torque as input. Considering most low-cost servos only support angle control, we hereby model the servo’s motion as a first-order model
@@ II-B3 Servo Model
where $\bm{\alpha}=\left[\alpha_{1},\alpha_{2},\dots,\alpha_{N_{p}}\right]^{T}$ denotes the vector containing all servo angles, and $\bm{\alpha}_{c}=\left[\alpha_{c1},\alpha_{c2},\dots,\alpha_{cN_{p}}\right]^{T}$ is the symbol for servo angle commands.
II-B4 Rigid-Body Model
Neglecting the aerodynamic drag during flying, the motion of the robot is caused only by gravitational force and rotor wrench. Using position ${{}^{W}\bm{p}}$, velocity ${}^{W}\bm{v}$, quaternion ${{}^{W}_{G}\bm{q}}$, and angular velocity ${{}^{G}\bm{\omega}}$ (angular velocity of $\left\{\mathcal{G}\right\}$ w.r.t $\left\{\mathcal{W}\right\}$ and expressed in $\left\{\mathcal{G}\right\}$) as states, a six-DoF rigid body dynamics can be established as follows
${{}^{W}\dot{\bm{p}}}$ $={{}^{W}\bm{v}},$ (6a)
${{}^{W}\dot{\bm{v}}}$ $=\left({{}^{W}_{G}\bm{R}(\bm{q})}\ {{}^{G}\bm{f}_{u}}+{{}^{W}\bm{f}_{d}}\right)/{m}+{{}^{W}\bm{{g}}},$ (6b)
${{}^{W}_{G}\dot{\bm{q}}}$ $=\frac{1}{2}\ {{}^{W}_{G}\bm{q}}\circ\mathcal{H}(^{G}\bm{\omega}),$ (6c)
${{}^{G}\dot{\bm{\omega}}}$ $=\bm{{I}}^{-1}\left(-{{}^{G}\bm{\omega}}\times\left({\bm{{I}}}\ {{}^{G}\bm{\omega}}\right)+{{}^{G}\bm{\tau}_{u}}+{{}^{G}\bm{\tau}_{d}}\right),$ (6d)
where $\bm{R}(\bm{q})$ denotes the rotation matrix converted from a quaternion, $\circ$ refers to quaternion multiplication, $\mathcal{H}(\cdot)$ means homogenizing a 3D vector $\mathcal{H}(\bm{p}):=[0,\bm{p}]^{T}$, ${{}^{G}\bm{f}_{u}}$ and ${{}^{G}\bm{\tau}_{u}}$ are expressed in (4a) and (4b), ${{}^{W}\bm{f}_{d}}$ and ${{}^{G}\bm{\tau}_{d}}$ are the force and torque caused by disturbances, as well as $m$, ${{}^{W}\bm{{g}}}=[0,0,{-g}]^{T}$, and $\bm{I}{\color[rgb]{0,0,0}=\texttt{diag}({I}_{xx},{I}_{yy},{I}_{zz})}$ are the mass, the gravity vector, and the inertia matrix, respectively.
The whole model established above is utilized for control and simulation. Note that this model is created for multirotors, and the tiltable-quadrotor is the special case when $N_{p}=4$.